1.    Solve cos(x)=2x, using fixed point iteration to 5 decimal places             

Do 5 iterations.    Solve , using fixed point iteration to 5 decimal places       



1
Expert's answer
2021-09-07T15:36:14-0400

Given that "cos(x)=2x"

Applying fixed iteration method:

"x=\frac{cos(x)}{2}\\
\therefore x_{n+1}=\frac{cos(x_n)}{2}\\"

Take "x_0=1\\"

So,

"x_1=\frac{cos \ 1}{2}=0.27015\\
x_2=\frac{cos(x_1)}{2}=\frac{cos(0.27015)}{2}=0.48186\\
x_3=\frac{cos(x_2)}{2}=\frac{cos(0.48186)}{2}=0.44306\\
x_4=\frac{cos(x_3)}{2}=\frac{cos(0.44306)}{2}=0.45172\\
x_5=\frac{cos(x_4)}{2}=\frac{cos(0.45172)}{2}=0.44984\\"